OpenStudy (oswaldmurphy):

Choose the equation below whose axis of symmetry is x = 0. y = x2 + 2x y = x2 − 16x + 58 y = x2 + 2 y = x2 − 4x + 2

OpenStudy (oswaldmurphy):

@welshfella @ganeshie8

OpenStudy (oswaldmurphy):

Ok, so I change all Y to 0.

OpenStudy (welshfella):

oh sorry x = 0 is the y axis

OpenStudy (oswaldmurphy):

Ok. Make x = 0. What do I do after that?

OpenStudy (welshfella):

so you need to chooses a function which will give you 2 values +/- a for one value of y.

OpenStudy (oswaldmurphy):

I thought it was x = -b/2a

OpenStudy (imstuck):

In my opinion it is easier to just know the translations of these types of graphs. They all move up and down or side to side in the same way if you know how to "read" the equation.

OpenStudy (welshfella):

yeah - that another way of attempting a solution . b/2a will give you the axis of symmetry for a parabola

OpenStudy (oswaldmurphy):

So is the value for C 2?

OpenStudy (welshfella):

yes @IMStuck

OpenStudy (welshfella):

No

OpenStudy (welshfella):

try looking at the transformations of the graphs Hint: what does the graph of y = x^2 look like?

OpenStudy (oswaldmurphy):

http://prntscr.com/c1zowp

OpenStudy (welshfella):

right What is that symmetrical about/

OpenStudy (oswaldmurphy):

it splits it down the middle at the 0

OpenStudy (welshfella):

yes so its symmetrical about x = 0 ( that is the y-axis)

OpenStudy (welshfella):

And y = x^2 + 2 is same shaped graph moved 2 units up the y axis.

OpenStudy (welshfella):

so C is your answer

OpenStudy (welshfella):

applying the -b/2a thing to 2x^2 + 2 you need to write it in the form y = 2x^2 + 0x + 2 so -b/2a = 0/ 4 = 0 so the axis of sym is x = 0